# Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments. The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”

### v2

@@ 3.1 Revenue Maximization
The seller can obtain at least half the maximum revenue by selling one level of positional good.
{proofsketch}
Consider an auxiliary screening problem of selling indivisible ordinary goods (e.g., apples), where a posted price $p^{*}=\operatorname*{arg\,max}_{p}p(1-F(p))$ is optimal. I show that (i) by selling one level of positional good, the seller can obtain at least $\frac{1}{2}\max p(1-F(p))$ because those who buy the good get at least a status level of $s=1/2$; and (ii) the maximum revenue from selling positional goods is strictly below $\max p(1-F(p))$ because of positional externalities. (See the complete proof in Appendix B.)
@@
The auxiliary screening problem considered in the proof, which is equivalent to selling a continuum of indivisible ordinary goods to a continuum of buyers, highlights the distinction between selling ordinary goods and positional goods. In the screening problem, selling an item to one consumer has no externalities on others. By contrast, negative externalities are present when selling positional goods—the seller cannot offer the highest status to multiple buyers at face value. Thus, the seller of positional goods is subject to an additional constraint, resulting in strictly less revenue.
@@
Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$. Figure 2 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$. Assume the intrinsic value $v(\theta)=0$ for simplicity.181818 A positive intrinsic value $v(\theta)\geq 0$ will increase $\frac{R_{1}}{\max R}$ by increasing both $R_{1}$ and $\max R$. In Figure 2, (i) the blue area in the left panel represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. (ii) The red area in the right panel represents the revenue from selling one level of positional good at price $F(\theta_{0}^{*})$—i.e., $\bar{s}(\theta)=\frac{1+F(\theta_{0}^{*})}{2}\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. 191919 To see (i) and (ii), the revenue under $s^{*}$ is $\int_{\theta_{0}^{*}}^{1}J(x)F(x)\mathop{}\!\mathrm{d}F(x)=-\int_{\tau^{*}_{0}}^{1}\tau\mathop{}\!\mathrm{d}R(\tau)=\tau^{*}_{0}R(\tau^{*}_{0})+\int_{\tau^{*}_{0}}^{1}R(\tau)\mathop{}\!\mathrm{d}\tau$. Analogously, the revenue under $\bar{s}$ is $\int_{\theta_{0}^{*}}^{1}J(x)\frac{1+F(\theta_{0}^{*})}{2}\mathop{}\!\mathrm{d}F(x)=\frac{1+\tau^{*}_{0}}{2}R(\tau^{*}_{0})$. Because the seller can optimize the price (and setting it at $F(\theta_{0}^{*})$ is feasible), the revenue from selling one level of positional good is at least the red area. (iii) The entire box in either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem. It is straightforward to see that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
@@
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, selling one level of positional good can obtain $91.9\%$ of the maximum revenue.
@@
If $F(\theta)=\theta$ on $[0,1]$, selling one level of positional good can obtain $92.4\%$ of the maximum revenue.
@@
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, selling one level of positional good can obtain
@@ 3.2 Welfare Implications
The consumer surplus (or agent welfare) is given by
@@
    Fixing exclusion $\theta_{0}$, 202020 The exclusion level is fixed when the price of the lowest-level good is fixed. In particular, when there is a free regular service, there is no exclusion (see Appendix A.2). the consumer surplus is decreasing in the number of positional good levels the seller offers.
@@
It is also worth noting that under DFR, if the distribution still satisfies Myerson’s regularity, the seller’s revenue and the consumer surplus are aligned—both are higher when the seller offers more levels of positional goods.
Consumer Surplus Maximization.
Relatedly, I study the optimal mechanism that maximizes consumer surplus. After rewriting the consumer surplus as
@@
it is straightforward to show that when negative transfers are allowed and subject to budget balance ($\operatorname{\mathbf{E}}[p(\theta)]=0$), the consumer surplus is maximized by full separation without exclusion. Intuitively, this is because assortative matching is efficient, and transfers can redistribute efficiency gains among consumers.212121 See also Gershkov and Schweinzer (2010). I thank Alex Gershkov for pointing this out.
@@
Negative prices are implausible in some applications. In particular, in the applications to organizational hierarchies and education, the price (i.e., effort) must be nonnegative. Because the nonnegative price constraint also pins down $U(\theta_{0})$ through (IR), the characterization of the consumer surplus-maximizing mechanism under the constraint is similar to Proposition 4.
@@
Under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$, the first (second) part still holds if $\gamma\in[1/2,1]$ ($\gamma\in[0,1/2]$). In particular, full separation is always optimal if $\gamma=0$ (Immorlica et al., 2015), while total pooling is always optimal if $\gamma=1$ (Frank, 1985b).
@@
The result shows that under the nonnegative price constraint, exclusion is detrimental to the consumer under either IFR or DFR. Under IFR, because it is optimal to pool participants, exclusion can be beneficial as it increases the status of the participants who would otherwise be pooled with the excluded consumers. The result shows that the marginal loss of the excluded consumers exceeds the inframarginal gain of participants, so exclusion is detrimental to the consumer. On the other hand, under DFR, exclusion has no benefits to participants because the optimal mechanism fully separate participants, so no exclusion is optimal.
The result implies that under IFR, the consumer-optimal regulation of positional goods without using subsidies is to restrict the seller to offer one level of positional good and serve all consumers (e.g., by providing a free regular boarding service).
The results also have several implications in organizational hierarchies and education, where the effort (price) has to be positive. On the one hand, if the type distribution satisfies the IFR property (i.e, has a thin tail), which is satisfied by many common distributions, completely eliminating status competitions (e.g., educational disarmament) is optimal for agent welfare. In other words, agents pay nothing and are pooled at the same status. In the education application, this implies complete educational disarmament—for example, random assignment in allocating slots at sought-after schools. On the other hand, if there are a few very high types (i.e., a heavy tail), such as the Pareto distribution, intense competition or meritocracy is optimal even from the agent’s perspective.
@@ 3.3 Social Planner
Assume that the social planner maximizes a weighted sum of the seller’s revenue and consumer surplus, so social welfare is $S=\int_{0}^{\bar{\theta}}\lambda p(\theta)+U(\theta)\mathop{}\!\mathrm{d}F(\theta)$, where $\lambda\geq 0$. In the monopolist provision problem where $p$ is transferred to the seller, $\lambda$ captures the weight of the seller’s revenue relative to the consumer surplus. When the price is allowed to be negative, I assume $\lambda\geq 1$, because otherwise the social planner would set $p(\theta)=-\infty$ to transfer all money from the seller to consumers.
In the education or the organization application, where $p$ is the agent’s nonnegative effort (so utilities are nontransferable), $\lambda$ captures the extent to which effort is socially productive. Agent effort is completely unproductive when $\lambda=0$ and productive when $\lambda=1$. When $\lambda>1$, the social benefit of effort is higher than the private benefit.
Denote $J_{\lambda}(\theta)=\lambda\theta-(\lambda-1)\frac{1-F(\theta)}{f(\theta)}$. By previous results, social welfare is
$S=\int_{\theta_{0}}^{\bar{\theta}}J_{\lambda}(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{0}}^{\bar{\theta}}\left(\lambda v(\theta)-(\lambda v^{\prime}(\theta)-1)\frac{1-F(\theta)}{f(\theta)}-(\lambda-1)U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta).$ (13)
@@
The social welfare-maximizing mechanism fully separates the participants and potentially exclude some types
@@ 4.1 Intrinsic Quality
In this subsection, I consider the case where the positional good has an intrinsic quality ${q}$, produced at a cost $c({q})$, which determined its intrinsic value $v({q},\theta)=\theta q$ (à la Mussa and Rosen (1978)). Assume $c({q})$ is strictly increasing and strictly convex. Denote by $Q(\theta)=\operatorname{\mathbf{E}}_{\omega}[\chi(\theta,\omega)]$ the interim allocation. Thus, the buyer’s utility is $U(\theta)=\theta(s(\theta)+Q(\theta))-p(\theta)$.
@@
Assume $J(\theta)$ is increasing. The revenue-maximizing mechanism excludes types below $\theta^{*}_{0}=J^{-1}(0)$ and offers $\chi^{*}(\theta)=c^{\prime-1}(J(\theta))$ to $\theta\geq\theta^{*}_{0}$.
@@ 4.2 Deliberately Lowering the Status
In some applications, the principal can deliberately lower the status. In the priority service example (Example 2.1), the seller can prolong the waiting time $t=1-s$ by pausing service or procrastinating. Thus, the seller can create negative status $s<0$ or excessive waiting time $t>1$. 222222 For example, in the U.S., OPT processing time is excessively long for the regular service but takes only 30 days for the premium service.
In these cases, the range of status allocation $s$ becomes $[-M,1]$, where $M$ is a sufficiently large bound on the negative status (or the maximum excessive waiting time). Incentive compatibility still requires $s$ to be increasing, and feasibility requires $s\leq\hat{s}$ for some $\hat{s}\in\operatorname{\text{MPS}}_{w}(F)$. By lowering the status, the seller can offer negative status $s<0$ to low types and reduces their payoffs. Therefore, negative status (or excessive waiting time) serves as a substitute for exclusion.
@@
Under IFR, allowing the seller to lower the status increases the seller’s revenue and decreases the consumer surplus. 232323 It strictly increases the seller’s revenue if $v$ is nonlinear (i.e., strictly concave or $v(0)>0$) and strictly decreases the consumer surplus if $v$ is nonlinear and $f(0)<\infty$.
@@
The allocation $s^{*}(\theta)=-v^{\prime}(\theta)$ is increasing because $v^{\prime\prime}(\theta)\leq 0$ and feasible because it can be induced by randomization.
Example 4.1 (Excessive Waiting Time).
Consider the payoff $u(p,s,\theta)=\theta s-p+v(\theta)$ ($=\tilde{v}(\theta)-\theta t-p$), where $t=1-s$ is the waiting time. Assume $\theta\sim\text{Unif }[0,1]$ and that $v(\theta)=v_{0}+\alpha\theta$, where $v_{0},\alpha>0$. Then, the revenue-maximizing mechanism is
@@
The insight that negative status serves as a substitute for exclusion also applies to the case where exclusion is impossible as the lowest-level good or service is free (see Appendix A). In this case, negative status or excessive waiting time diminishes the value of the free service and effectively plays the role of exclusion, thereby increasing the seller’s revenue. In particular, by offering a priority service and a free regular service, the seller extracts more revenue as she increases the waiting time for the free regular service.

### v3

@@ 3.1.1 Approximation
The seller can obtain at least half the maximum revenue by selling a single good.
Proof sketch.
Consider an auxiliary screening problem of selling indivisible ordinary goods (e.g., apples), where a posted price $p^{*}=\operatorname*{arg\,max}_{p}p(1-F(p))$ is optimal. I show that (i) by selling a single good, the seller can obtain at least $\frac{1}{2}\max p(1-F(p))$ because those who buy the good get at least a status level of $s=1/2$; and (ii) the maximum revenue from selling positional goods is strictly below $\max p(1-F(p))$ because of positional externalities. (See the complete proof in Appendix B.) ∎
@@
The auxiliary screening problem considered in the proof, which is equivalent to selling a continuum of indivisible ordinary goods to a continuum of buyers, highlights the distinction between selling ordinary goods and positional goods. In the screening problem, selling an item to one consumer has no externalities on others. By contrast, when selling positional goods, because consumers care about their relative positions, negative externalities are present—the seller cannot offer the highest status to multiple buyers without reducing its value. Thus, the seller of positional goods is subject to an additional feasibility constraint, resulting in a strictly lower revenue.
@@
Figure 2 plots the revenue curve $R(\tau)=(1-\tau)F^{-1}(\tau)$ in the quantile space $\tau=F(\theta)$. Assume the intrinsic value $v(\theta)=0$ for simplicity. [^18] A positive intrinsic value $v(\theta)\geq 0$ will increase the ratio $\frac{R_{1}}{\max R}$, as it increases $R_{1}$ by at least the same amount as $\max R$. Then, the blue area in Figure 2 represents the revenue from the revenue-maximizing $s^{*}(\theta)=F(\theta)\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. To see this, note that $R^{\prime}(\tau)=-J(\tau)$, so using a change of variable $\tau=F(\theta)$ (and $\tau_{0}^{*}=F(\theta_{0}^{*})$) and integration by parts, the revenue under $s^{*}(\theta)$ is
$\int_{\theta_{0}^{*}}^{1}J(\theta)F(\theta)\mathop{}\!\mathrm{d}F(\theta)=-\int_{\tau^{*}_{0}}^{1}\tau\mathop{}\!\mathrm{d}R(\tau)=\tau^{*}_{0}R(\tau^{*}_{0})+\int_{\tau^{*}_{0}}^{1}R(\tau)\mathop{}\!\mathrm{d}\tau,$
which is the area of the blue region.
@@
Moreover, the red trapezoid in the right panel represents the revenue from selling a single good at price $F(\theta_{0}^{*})$—i.e., $\bar{s}(\theta)=\frac{1+F(\theta_{0}^{*})}{2}\cdot\mathbf{1}[\theta\geq\theta_{0}^{*}]$. This is because the revenue under $\bar{s}$ is
$\int_{\theta_{0}^{*}}^{1}J(\theta)\frac{1+F(\theta_{0}^{*})}{2}\mathop{}\!\mathrm{d}F(\theta)=\frac{1+\tau^{*}_{0}}{2}R(\tau^{*}_{0}),$
which is the area of the red trapezoid in the right panel. Because the seller can optimize the price (and setting it at $F(\theta_{0}^{*})$ is feasible), the revenue from selling a single good is at least the red area.
Finally, the entire box, which has an area of $R(\tau^{*}_{0})$, represents the maximum revenue from selling nonpositional goods in the auxiliary problem. It is straightforward to see that the red area is larger than half of the entire box and thus strictly larger than half of the blue area.
@@
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, selling a single good can obtain $91.9\%$ of the maximum revenue.
@@
If $F(\theta)=\theta$ on $[0,1]$, selling a single good can obtain $92.4\%$ of the maximum revenue.
@@
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, selling a single good can obtain
@@ 3.2 Welfare Implications
The consumer surplus (or agents’ welfare) is given by
@@
    Fixing exclusion $\theta_{0}$,  [^19] The exclusion level is fixed when the price of the lowest-level good is fixed. In particular, when there is a free regular service, there is no exclusion (see Appendix A.2). the consumer surplus is decreasing in the number of positional good levels the seller offers.
@@
It is also worth noting that if the distribution satisfies Myerson’s regularity but has DFR (i.e., a heavy tail), the seller’s revenue and the consumer surplus are aligned—both are higher when the seller offers more levels of positional goods. Intuitively, offering more levels of positional goods benefits high types and hurts low types, so it benefits agents in aggregate when there are relatively more high types.
3.2.1 Consumer Surplus Maximization
Consider next the optimal mechanism that maximizes consumer surplus. After rewriting the consumer surplus as
@@ 3.2.1 Consumer Surplus Maximization
it is straightforward to show that when negative transfers are allowed and subject to budget balance ($\operatorname{\mathbf{E}}[p(\theta)]=0$), the consumer surplus is maximized by full separation without exclusion. Intuitively, this is because assortative matching is efficient, and transfers can redistribute efficiency gains among consumers and induce truth-telling. [^20] See also Gershkov and Schweinzer (2010). I thank Alex Gershkov for pointing this out.
@@
The above result relies on the assumption that negative transfers are allowed, which is implausible in some applications. In particular, in the applications to organizational hierarchies and education, the price (i.e., effort) must be nonnegative. Because the nonnegative price constraint also pins down $U(\theta_{0})$ through (IR), the characterization of the consumer surplus-maximizing mechanism under the constraint is similar to Proposition 4.
@@
Under the specification $\tilde{S}(q,G)=\gamma G({q})+(1-\gamma)G^{-}({q})$, the first (second) part still holds if $\gamma\in[1/2,1]$ ($\gamma\in[0,1/2]$). In particular, full separation is always optimal if $\gamma=0$ (Immorlica, Stoddard and Syrgkanis, 2015), while total pooling is always optimal if $\gamma=1$ (Frank, 1985b).
@@
The result shows that under the nonnegative price constraint, exclusion is detrimental to the consumer under either IFR or DFR. Under IFR, because it is optimal to pool participants, exclusion can be beneficial as it increases the status of the participants who would otherwise be pooled with the excluded consumers. The result shows that the marginal loss of the excluded consumers exceeds the inframarginal gain of participants, so exclusion is detrimental to the consumer. On the other hand, under DFR, exclusion has no benefits to participants because the optimal mechanism fully separates participants, so no exclusion is optimal.
The result implies that under IFR, the consumer-optimal regulation of positional goods without using subsidies is to restrict the seller to offer a single good and serve all consumers (e.g., by providing a free regular boarding service).
The results also have several implications in organizational hierarchies and education, where the effort (price) has to be positive. On the one hand, if the type distribution satisfies the IFR property (i.e, has a thin tail), completely eliminating status competitions (e.g., educational disarmament) is optimal for agents’ welfare. In the education application, this implies complete educational disarmament—for example, random assignment in allocating slots at sought-after schools. On the other hand, if there are a few very high types (i.e., a heavy tail), such as the Pareto distribution, meritocracy is optimal for agents’ welfare. Intuitively, low types benefit from soft competition, while high types prefer intense competition; hence, on average, more (less) intense competition is optimal for agents’ welfare when the population contains relatively more high (low) types.
@@ 3.3 Social Planner
Assume that the social planner maximizes a weighted sum of the seller’s revenue and consumer surplus, so social welfare is $W_{S}=\int_{0}^{\bar{\theta}}\lambda p(\theta)+U(\theta)\mathop{}\!\mathrm{d}F(\theta)$, where $\lambda\geq 0$. In the monopolist provision problem where $p$ is transferred to the seller, $\lambda$ captures the weight of the seller’s revenue relative to the consumer surplus. When the price is allowed to be negative, I assume $\lambda\geq 1$, because otherwise the social planner would transfer all money from the seller to consumers.
In the education or the organization application, where $p$ is the agent’s nonnegative effort (so utilities are nontransferable), $\lambda$ captures the extent to which effort is socially productive. agents’ effort is completely unproductive when $\lambda=0$ and productive when $\lambda=1$. When $\lambda>1$, the social benefit of effort is higher than the private benefit.
Define $J_{\lambda}(\theta)=\lambda\theta-(\lambda-1)\frac{1-F(\theta)}{f(\theta)}$. The expected social welfare is
$W_{S}=\int_{\theta_{0}}^{\bar{\theta}}J_{\lambda}(\theta)s(\theta)\mathop{}\!\mathrm{d}F(\theta)+\int_{\theta_{0}}^{\bar{\theta}}\left(\lambda v(\theta)-(\lambda v^{\prime}(\theta)-1)\frac{1-F(\theta)}{f(\theta)}-(\lambda-1)U(\theta_{0})\right)\mathop{}\!\mathrm{d}F(\theta).$ (13)
@@
The social welfare-maximizing mechanism fully separates the participants and potentially excludes some types
@@ 4.1 Intrinsic Quality
I have maintained the assumption that the positional good has no intrinsic value to abstract from quality differentiation. However, in many applications, a higher-level positional good (e.g., a larger car or higher boarding class) also comes with a higher intrinsic value. Now assume the positional good q{q} has an intrinsic quality produced at a cost c(q)c({q}), which determines its intrinsic value $v({q},\theta)=\theta q$, à la Mussa and Rosen (1978). Assume c(q)c({q}) is strictly increasing and strictly convex. Denote by $Q(\theta)=\operatorname{\mathbf{E}}_{\omega}[\chi(\theta,\omega)]$ the interim allocation. Thus, the buyer’s utility is $U(\theta)=\theta(s(\theta)+Q(\theta))-p(\theta)$.
@@
Assume $J(\theta)$ is increasing. The revenue-maximizing mechanism excludes types below $\theta^{*}_{0}=J^{-1}(0)$ and offers $Q^{*}(\theta)=c^{\prime-1}(J(\theta))$ to $\theta\geq\theta^{*}_{0}$.
@@
4.2 Convex or Concave Status
When positional externalities arise from interpersonal comparison, a concave or convex specification of status may be plausible. Formally, assume the status function is
$\tilde{S}({q},G)=\phi(S({q},G))=\phi\left(\frac{G^{-}({q})+G({q})}{2}\right),$ (15)
where $\phi(\cdot)$ is a strictly increasing function. Under this specification, an increasing interim status $\tilde{s}$ is feasible if and only if $\tilde{s}\in\operatorname{conv}\{\sigma=\phi\circ s\mid s\in\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)\}$. Because the maximization problems remain linear in $\tilde{s}$, the optimal $\tilde{s}^{*}$ is an extreme point of the feasible set, which satisfies $\tilde{s}^{*}=\phi\circ s^{*}$ for some $s^{*}\in\operatorname{ext}\operatorname{\text{MPS}}_{w}(F)$. Hence, the following proposition shows that the previous results are robust to concave or convex status, although they may require stronger regularity conditions.
Proposition 8.
(i) If $\phi$ is strictly increasing and convex, Propositions 2, 4, 5, and 6 still hold (with a different exclusion level) after replacing the IFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is decreasing) with
$\frac{1-F(\theta)}{f(\theta)}\phi^{\prime}(F(\theta))\mbox{ is decreasing.}$
(ii) If $\phi$ is strictly increasing and concave, Propositions 2, 4, 5, and 6 still hold (with a different exclusion level) after replacing
  • Myerson’s regularity (i.e., $J(\theta)$ is increasing) with
    $J(\theta)\phi^{\prime}(F(\theta))\mbox{ is increasing;}$
  • the DFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is increasing) with
    $\frac{1-F(\theta)}{f(\theta)}\phi^{\prime}(F(\theta))\mbox{ is increasing;}$
  • $J_{\lambda}(\theta)$ is increasing with
    $J_{\lambda}(\theta)\phi^{\prime}(F(\theta))\mbox{ is increasing.}$
In cases where conditions are unchanged (e.g., in revenue maximization if $\phi$ is convex), the result follows from the Fan-Lorentz Theorem. When conditions are stronger, the proof uses the optimal control method.
@@ 4.3 Deliberately Lowering the Status
In some applications, the principal can deliberately lower the status. In the priority service example (Example 1), the seller can prolong the waiting time t=1st=1-s by pausing service or procrastinating. Thus, the seller can create excessive waiting time $t>1$ or essentially negative status $s<0$.
In these cases, the range of status allocation $s$ becomes [M,1][-M,1], where $M>0$ is a sufficiently large bound on the negative status (or the maximum excessive waiting time). Incentive compatibility still requires $s$ to be increasing, and feasibility requires $s\leq\hat{s}$ for some $\hat{s}\in\operatorname{\text{MPS}}_{w}(F)$. By lowering the status, the seller can offer negative status $s<0$ to low types and reduce their payoffs. Therefore, negative status (or excessive waiting time) serves as a substitute for exclusion.
@@
Under IFR, allowing the seller to lower the status increases the seller’s revenue and decreases the consumer surplus.  [^21] It strictly increases the seller’s revenue if $v$ is nonlinear (i.e., strictly concave or $v(0)>0$) and strictly decreases the consumer surplus if $v$ is nonlinear and $f(0)<\infty$.
@@
The allocation $s^{*}(\theta)=-v^{\prime}(\theta)$ is increasing because $v^{\prime\prime}(\theta)\leq 0$ and feasible because it can be induced by randomization between pooling and separation (after lowering the status by a constant).
Example 3 (Excessive Waiting Time).
Consider the payoff $u(p,s,\theta)=\theta s-p+v(\theta)$ ($\tilde{v}(\theta)-\theta t-p$), where t=1st=1-s is the waiting time. Assume $\theta\sim\text{Unif }[0,1]$ and that $v(\theta)=v_{0}+\alpha\theta$, where $v_{0},\alpha>0$. Then, the revenue-maximizing mechanism is
@@
The insight that negative status serves as a substitute for exclusion also applies to the case where exclusion is impossible, as the lowest-level good or service is free (see Appendix A). In this case, negative status or excessive waiting time diminishes the value of the free service and effectively plays the role of exclusion, thereby increasing the seller’s revenue. In particular, by offering a priority service and a free regular service, the seller extracts more revenue as she increases the waiting time for the free regular service.
4.4 Signaling
Consider an alternative specification where the status arises from signaling concerns. For example, consumers buy conspicuous goods $q$ to obtain social status with a value of $\operatorname{\mathbf{E}}[\theta\mid q]$ (Rayo, 2013). Under this specification, the status value induced by a positional good allocation $\chi(\theta)$ is given by $S({q},G_{\chi})=\operatorname{\mathbf{E}}[\theta\mid{\chi(\theta)=q}]$. Moreover, even if the consumer does not buy the good (i.e., $q=0$), he can still obtain some status value $\underline{s}=\operatorname{\mathbf{E}}[\theta\mid q=0]$, instead of zero in the main specification. Therefore, exclusion is impossible because the consumer can get the lowest status $\underline{s}$ for free.
Analogous to Theorem 1 (without exclusion), an increasing interim status $\tilde{s}$ is feasible if and only if $s\in\operatorname{\text{MPS}}(\theta)$ in the quantile space. Therefore, the main results continue to hold except that exclusion is impossible (see also Appendix A). In particular, the revenue-maximizing mechanism fully separates all types if and only if $J(\theta)$ is increasing, while selling more levels reduces the consumer surplus if $F$ satisfies the IFR property.